When rolling a pair of standard dice, what is the probability of rolling a sum that is less than 10?
Found 2 solutions by ikleyn, ewatrrr:
Answer by ikleyn(52788)

.

If you want to learn this subject and this class of problems,  look into the lesson
    - Rolling a pair of fair dice
in this site.  You will find there many similar solved problems.

Also,  you have this free of charge online textbook in ALGEBRA-II in this site
    - ALGEBRA-II - YOUR ONLINE TEXTBOOK.

The referred lesson is the part of this online textbook under the topic  "Solved problems on Probability".


Save the link to this textbook together with its description

Free of charge online textbook in ALGEBRA-II
https://www.algebra.com/algebra/homework/complex/ALGEBRA-II-YOUR-ONLINE-TEXTBOOK.lesson

into your archive and use when it is needed.



Answer by ewatrrr(24785)

 
Hi,
 When rolling a pair of standard dice.  36 ways it can happen 
what is the probability of rolling a sum that is less than 10? 
 P(sum < 10) = 1 - P (rolling sum ≥ 10) = 1- 6/36 = 30/36 = 5/6  
  P(sum < 10) = 5/6   0r .8333  
know if it should be left as a fraction 0r decimal (rounded off as directed).
Wish You the Best in your Studies.
  -----------------
+__1__2__3__4__5__6
1|_2__3__4__5__6__7
2|_3__4__5__6__7__8
3|_4__5__6__7__8__9
4|_5__6__7__8__9__10
5|_6__7__8__9__10_11
6|_7__8__9_10__11_12